OpenStudy (anonymous):

How would you write an equation of the parabola with a directrix y = -1 and vertex (0,0)?

OpenStudy (anonymous):

as the vertex is the origin there's no constant term

OpenStudy (amistre64):

the direct-x is the line that is equal distance from the curve of the graph to the focus of the parabola. If the directix is at y=-1 then the focus is at y=1. I am guessing that the equation is (x^2), but I recall the equation of a parabola taking a slightly different form when you write it out geometrically.

OpenStudy (anonymous):

given the focus point its easier to find the equation of this parabola, as the equation is (x-h)^2=4p(y-k) where (h,k)=(0,0)

OpenStudy (anonymous):

Thanks, got it.

OpenStudy (anonymous):

p is the distance between vertex and directrix..try to find the value using this

OpenStudy (anonymous):

so, whats the equation?

OpenStudy (anonymous):

y = 1/4x^2

OpenStudy (anonymous):

good